# Beyond Incompatibility: Trade-offs between Mutually Exclusive Fairness Criteria in Machine Learning and Law

### v5

@@ 7 Conclusion
Our second main contribution is to map these technical interpolation capacities to legal provisions requiring nuanced trade-offs between fairness, performance and other criteria relating to the protection of fundamental rights. Operationalizations of the often vague legal terms are crucial for compliance and to deploy AI systems in ways that are both safe and legal. We offer novel interpretations of the recently enacted EU AI Act and the Digital Markets Act against this technical background, and show howFAIM may help navigate these emerging, new frontiers at the intersection of law and computer science.
Hence, we have also stressed that any choice of $\theta$ values must, of course, remain within the confines of the law. We discussed how FAIM may be harnessed to comply with competing legal non-discrimination obligations in three concrete settings corresponding to our experiments: credit scoring, bail decisions, and emerging notions of fairness in e-commerce rankings. The discussion shows that our model can handle both traditional legal fields and novel obligations under the recently enacted AI Act and the DMA. To our knowledge, our paper is the first to explore these new instruments from both a technical and a legal perspective.

### v6

@@ 6 Discussion and Limitations
6.1 Choosing Fairness Weights
In the three discussed use cases, decision makers may place differing emphasis on various fairness criteria depending on which undesirable error type is most relevant. In recidivism prediction, balancing both error types might appear important, but judges or lawmakers may weight false positive predictions more heavily if unwarranted detentions are considered a particularly severe intrusion into personal freedom. In credit scoring, decision makers may choose to weight false negative rates more highly in high-stakes loan scenarios, while low-stakes loans may justify placing greater emphasis on false positives to avoid unnecessarily denying credit access. In e-commerce rankings, decision makers may prefer to focus on equalizing false positives if both accuracy and fairness in the top-ranked products are critical, or choose to emphasize false negatives if it is necessary to ensure that products are not unfairly pushed to low-visibility positions.
More generally, the selection of weights for fairness criteria in different practical applications will vary, as the real-world consequences of imbalance differ significantly across contexts. Decision makers need to consider not only what each error type entails in theory, but also what it means in the given environment.
It is often useful to start with a baseline weighting of 1/3 each and then adjust these weights based on an analysis of the stakeholders’ priorities, the severity of potential harms, and the anticipated legal and ethical constraints. In settings where reputation management is critical, decision makers may give equal weighting to all three criteria by setting them to 0.333 each, ensuring that no single measure dominates. In scenarios where a specific error type may lead to legal liability, that criterion could receive a higher weight such as 0.6 or 0.7, with the remaining 0.2 or 0.3 split evenly between the other two criteria. In the following paragraphs, we discuss briefly considerations for hiring decisions, college admissions, and medical recommendations.
Hiring decisions: In this context, false negatives (qualified candidates not hired) and false positives (unqualified candidates hired) can both have significant consequences. However, the harm of unfairly denying opportunities likely outweighs the cost of some suboptimal hires, unless the position is for a unique leadership role. Therefore, we would suggest generally emphasizing equal false negative rates between demographic groups (e.g. by sex, race) with a weight of 0.5 for Criterion C, 0.3 for equalizing false positives (Criterion B), and 0.2 for the Criterion A (calibration). This allows a company to draw on a more diverse set of candidates and experiment with different protected groups whose potential might otherwise be overlooked. If, however, the role is of particular importance for the company and quite unique (e.g., C-level position), avoiding false positives seems particularly desirable and justified. Here, Criterion B could therefore be pushed to 0.6 or 0.7, with the remaining weights split between the other two.
College admissions: Similar to the general hiring scenario, the impact of a false negative (unfairly denying a student admission) is arguably greater than a false positive (admitting a less-qualified student). Students have a limited number of opportunities to attend top colleges. Conversely, admitting a few suboptimally qualified students (false positives) does not hurt colleges or universities much. Hence, equalizing false negative rates should be the priority, perhaps with a 0.6 weight on Criterion C, 0.2 on Criterion B, and 0.2 on overall calibration.
Medical diagnosis and treatment recommendations: When applying a fairness algorithm like FAIM to medical diagnosis and treatment recommendations, setting the appropriate weights for the different fairness criteria is of utmost importance given the high-stakes nature of healthcare decisions [86]. False negative and false positive predictions can both lead to serious consequences for patients, so carefully considering the potential impact of each type of error is essential. False negatives, where a patient with a serious condition is incorrectly given a clean bill of health, can result in delayed diagnosis and treatment. This is particularly concerning for conditions like cancer, where early detection and intervention are critical. If an algorithm has a higher false negative rate for certain demographic groups, such as older patients, it could lead to a disproportionate number of those individuals facing worse prognoses or even preventable deaths due to missed or delayed diagnoses. False positives, on the other hand, can cause undue psychological stress, unnecessary medical procedures, and increased healthcare costs. Being incorrectly told you may have a serious illness can take a significant emotional and mental toll, even if subsequent tests reveal the initial diagnosis was wrong. False positives may also lead to invasive procedures like biopsies or surgeries that carry physical risks and financial burdens for the patient.
The specific FAIM weights will depend on the characteristics of the medical condition in question and the relative risks associated with false negatives and false positives. For a life-threatening illness where a missed diagnosis could be deadly, a higher weight for Criterion C may be justified, such as 0.6, while Criterion B might be assigned a weight of 0.3, and overall calibration 0.1. For less severe conditions, where false positives pose a bigger problem, the weights could be adjusted accordingly.
Overall, aligning weight distributions with the underlying normative trade-offs relevant to the specific domain allows decision makers to create a more justified and context-sensitive approach to implementing fair classification schemes.
@@ 7 Conclusion
Our second main contribution is to map these technical interpolation capacities to legal provisions requiring nuanced trade-offs between fairness, performance and other criteria relating to the protection of fundamental rights. Operationalizations of the often vague legal terms are crucial for compliance and to deploy AI systems in ways that are both safe and legal. We offer novel interpretations of the recently enacted EU AI Act and the Digital Markets Act against this technical background, and show how FAIM may help navigate these emerging, new frontiers at the intersection of law and computer science. Guidance is offered on how the trade-off between different fairness metrics may be conducted in various high-stakes settings, including hiring, college admissions, and healthcare.
In doing so, we have also stressed that any choice of $\theta$ values must, of course, remain within the confines of the law. We discuss how FAIM may be harnessed to comply with competing legal non-discrimination obligations in three concrete settings corresponding to our experiments: credit scoring, bail decisions, and emerging notions of fairness in e-commerce rankings. The discussion shows that our model can handle both traditional legal fields and novel obligations under the recently enacted AI Act and the DMA. To our knowledge, our paper is the first to explore these new instruments from both a technical and a legal perspective.